# Physics-Informed Weakly Supervised Learning for Interatomic Potentials

### v1

@@ 4.2 Physics-Informed Spatial-Consistency Loss
This section introduces a physics-informed approach for generating weak labels based on the concept of conservative forces. Thus, we leverage that the energy difference between two points on the potential energy surface is independent of the path taken between them. We consider two paths from a reference point to the same target point, composed of three perturbation vectors in total. We estimate the potential energy at the target point via Eq. (2). An example of two paths is demonstrated in Fig. 1 (c). The figure relates the energy obtained when displacing atomic positions of the original configuration $\mathcal{S}$ (denoted by A in the figure) by $\delta\mathbf{r}$ (from configuration A to C) with the energy obtained through consecutive perturbations $\delta\mathbf{r}^{\prime}$ (from configuration A to B) and $\delta\mathbf{r}^{\prime\prime}$ (from configuration B to C).
For the first path, we directly predict the energy with an MLIP, i.e., $E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)$, which is related to the approximated energy at $\mathbf{r}+\delta\mathbf{r}$ using Eq. (3) through PITC loss. For the second path, we directly compute the energy $E\left(\mathcal{S}_{\delta\mathbf{r}^{\prime}};\bm{\theta}\right)$ for atomic positions displaced by $\delta\mathbf{r}^{\prime}$ and use it to approximate $E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)$ after applying the second perturbation vector $\delta\mathbf{r}^{\prime\prime}\equiv\delta\mathbf{r}-\delta\mathbf{r}^{\prime}$. The physics-informed spatial consistency (PISC) loss can be defined as
$\begin{split}L_{\mathrm{PISC}}\left(\mathcal{S};\bm{\theta}\right)&=\ell\Big{(}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right),E\left(\mathcal{S}_{\delta\mathbf{r}^{\prime}};\bm{\theta}\right)-\sum_{i=1}^{N_{\mathrm{at}}}\langle\delta\mathbf{r}_{i}^{\prime\prime},\mathbf{F}_{i}\left(\mathcal{S}_{\delta\mathbf{r}^{\prime}};\bm{\theta}\right)\rangle\Big{)},\end{split}$ (4)
where $\delta\mathbf{r}$ is randomly-sampled or determined adversarially; see Section 4.4. After joint training of PITC and PISC losses, the three different estimations at $\mathcal{S}_{\delta\mathbf{r}}$ become spatially consistent. Note that our conservative forces-based approach is not limited to relations between two perturbation paths or three perturbation vectors. We discuss several other possible configurations in Section D.3.
@@ 4.3 Combined Physics-Informed Weakly Supervised Loss
Together with the usual MLIP loss function given in Eq. (1), the overall objective, which we refer to as the PIWSL loss, can be written as
$\operatorname*{arg\,min}_{\bm{\theta}}\tilde{\mathcal{L}}\left(\mathcal{D};\bm{\theta}\right)=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{\mathcal{S}\in\mathcal{D}}\left(L\left(\mathcal{S};\bm{\theta}\right)+C_{\mathrm{PITC}}L_{\mathrm{PITC}}\left(\mathcal{S};\bm{\theta}\right)+C_{\mathrm{PISC}}L_{\mathrm{PISC}}\left(\mathcal{S};\bm{\theta}\right)\right),$ (5)
@@ 4.4 Determining Perturbation Directions and Magnitudes
The effectiveness of the proposed approach depends on appropriate choices of the perturbation vectors $\delta\mathbf{r}$. Therefore, we introduce and justify various generation strategies for the perturbations $\delta\mathbf{r}$ used in Eq. (3) and Eq. (4). We can express any vector $\delta\mathbf{r}$ as $\delta\mathbf{r}\equiv\epsilon\mathbf{g}/\|\mathbf{g}\|_{2}$, where $\epsilon$ is the magnitude of $\delta\mathbf{r}$ and $\mathbf{g}/\|\mathbf{g}\|_{2}$ represents the direction of $\delta\mathbf{r}$. Physical constraints are considered when determining $\epsilon$. Specifically, the maximum length of the perturbation can be obtained from the validity of the Taylor expansion in Eq. (2), which, as discussed in Section 5.3, is typically given as at most 30% of the original bond length whose shortest example is the bond between carbon and hydrogen, about 1.09 Å; see also Fig. 2 (c) and (d). The specific values of $\epsilon$ chosen for our experiments are provided in Section B.1.
@@
Second, we compute an adversarial direction, as proposed elsewhere [48, 49], which involves defining it as the direction (the gradients) in which the loss error increases the most at the current atom coordinates $\mathbf{r}$ and for the current predicted energy. Assuming the norm of adversarial perturbation as $L_{2}$, the adversarial direction can be approximated by [49]
@@
where $L_{\rm dist}$ is a distance measure function to be maximized by adding $\delta\mathbf{r}_{\rm adv}$, with $\mathbf{y}^{\rm pred}$ and $\mathbf{y}^{\rm ref}$ being the ML model prediction and the reference values. Due to their computational efficiency, we mainly use Eq. (6) in our experiments. A quantitative comparison between the random and adversarial directions is provided in Section 5.5.
@@ 5 Experiments
We evaluate our method with an extensive set of experiments with the following objectives: (1) a comparison of the PIWSL method with existing baselines, (2) a detailed analysis of the impact of the PIWSL method using the aspirin molecule, (3) an exploration of the ability of PIWSL to improve energy and force predictions when force labels are inaccessible, (4) a comparison to prior work that used a weakly supervised approach, (5) an ablation study, and (6) a comparison between random and adversarial generation of the perturbation vector. In general, we focus on the data-scarce setting where the number of training samples is between $100$ and $1000$ since the computational cost of ab initio and first-principles approaches to generate large data sets can be prohibitive.
@@ 5.1 Models and Data Sets
We trained the following representative models that are provided in the Open Catalyst code base [9]: SchNet  [28], PaiNN [34], SpinConv [35], eSCN [36], and Equiformer v2 [37], covering MLIPs with a smaller (SchNet, SpinConv) and larger number of parameters (eSCN, Equiformer v2). Moreover, we also considered a parameter-efficient PaiNN model. Unless otherwise mentioned and except for SchNet, forces are directly predicted and not computed through the negative gradient of the energy. The results where forces are computed as negative energy gradients are analyzed in Section 5.4 and Section D.9. To evaluate the effect and dependency of the physics-informed weakly supervised approach in detail, we performed the training on various data sets: ANI-1x as a heterogeneous molecular data set [19], TiO2 as a data set for inorganic materials [26], the revised MD17 (rMD17) data set containing small molecules with sampled configurational spaces for each [46, 50, 51], and LMNTO as another material data set [17]; the results for rMD17 and LMNTO are provided in Section D.1. The detailed description of each data set is provided in Section B.3.
@@ 5.2 Benchmark Results
We compare models trained using the PIWSL loss (see Eq. (5)) with baseline models trained using the standard supervised loss only (see Eq. (1)). We also compare our approach to a recently proposed data augmentation method that incorporates the task of denoising random perturbations of the atomic coordinates into the learning objective (NoisyNode) [43]. More details on the setup are provided in Section B.1.
Heterogeneous Molecular Data Set (ANI-1x). The results provided in Table 1 show that our approach improves the baseline models’ performance in almost all cases. In particular, the error reduction for the predicted energies is often between $10$ and more than $50$. Interestingly, we observe an improved accuracy for potential energies and atomic forces because we include force prediction in PITC and PISC losses, different from the previous work [17]. In most cases, except for SchNet, the data augmentation method (NoisyNode) degrades the accuracy of the MLIPs because it does not incorporate the proper response of the energy and atomic forces to the perturbation of atomic positions.
@@
$N_{\mathrm{train}}=100$ $N_{\mathrm{train}}=1000$
@@
SchNeta E 17.21 $\pm$ 0.00 19.68 $\pm$ 0.00 17.08 $\pm$ 0.00 9.56 $\pm$ 0.00 9.60 $\pm$ 0.00 9.51 $\pm$ 0.00
F 2.84 $\pm$ 0.00 2.70 $\pm$ 0.00 2.83 $\pm$ 0.00 2.14 $\pm$ 0.00 2.15 $\pm$ 0.00 2.13 $\pm$ 0.00
PaiNNb E 14.41 $\pm$ 0.16 n/ab 13.95 $\pm$ 0.09 4.49 $\pm$ 0.15 n/ab 3.63 $\pm$ 0.20
F 1.59 $\pm$ 0.01 n/ab 1.56 $\pm$ 0.01 0.41 $\pm$ 0.02 n/ab 0.34 $\pm$ 0.01
SpinConv E 20.00 $\pm$ 0.42 18.76 $\pm$ 0.74 16.98 $\pm$ 0.99 4.17 $\pm$ 0.76 4.09 $\pm$ 0.65 2.50 $\pm$ 0.40
F 1.58 $\pm$ 0.03 1.53 $\pm$ 0.03 1.59 $\pm$ 0.03 0.65 $\pm$ 0.02 0.71 $\pm$ 0.16 0.58 $\pm$ 0.05
eSCN E 16.41 $\pm$ 1.10 20.92 $\pm$ 0.00 12.63 $\pm$ 0.78 3.31 $\pm$ 1.18 20.90 $\pm$ 0.01 1.40 $\pm$ 0.10
F 1.57 $\pm$ 0.04 1.66 $\pm$ 0.00 1.44 $\pm$ 0.03 0.46 $\pm$ 0.23 1.66 $\pm$ 0.00 0.21 $\pm$ 0.00
Equiformer E 18.21 $\pm$ 0.02 19.06 $\pm$ 0.02 13.93 $\pm$ 0.09 3.67 $\pm$ 0.03 18.75 $\pm$ 0.05 1.82 $\pm$ 0.34
F 1.56 $\pm$ 0.01 1.64 $\pm$ 0.00 1.51 $\pm$ 0.19 0.17 $\pm$ 0.01 1.58 $\pm$ 0.00 0.17 $\pm$ 0.01
@@
a We used a larger batch size of $32$ for SchNet since we obtained extremely high errors for the batch size of $4$. A more detailed discussion of the experimental results for SchNet is provided in Section D.1.
@@
b Because of a numerical instability of PaiNN when perturbing atomic coordinates, the cutoff radius is reduced from 12 Å to 5 Å in this experiment. Predicted values become n/a when perturbations are added to atomic configurations.
@@
Figure 2: (a, b) Relative performance gains for MLIPs trained with PIWSL compared to those trained without it and (c, d) Potential energy profiles for a C–H bond of the aspirin molecule. Relative performance gains are evaluated for (a) energy (E-) and (b) force (F-) RMSEs. These results are presented for the ANI-1x data set. Potential energy profiles for a C–H bond of the aspirin molecule are presented for models trained using (c) 100 and (d) 200 configurations. The red and blue arrows indicate the direction from the original structure ($E(\mathcal{S};\bm{\theta})$) to the perturbed one ($E(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta})$), as defined by Eq. (2), for the baseline and PIWSL model predictions, respectively.
Table 3: Results for models trained on the ANI-1x data set without reference forces. All models are trained using 1000 training samples. FB refers to the setting where the force branch estimates the force, and GB denotes the setting where the force is estimated by the gradient of the potential energy with respect to the atomic coordinates.
Model Case Baseline PIWSL
PaiNN FB E 42.36 $\pm$ 0.30 25.42 $\pm$ 0.72
F 24.25 $\pm$ 0.00 20.54 $\pm$ 0.08
GF E 41.83 $\pm$ 1.81 29.71 $\pm$ 0.55
F 83.36 $\pm$ 2.85 24.02 $\pm$ 0.95
Equiformer FB E 43.14 $\pm$ 0.86 29.48 $\pm$ 0.51
F 24.25 $\pm$ 0.00 21.99 $\pm$ 0.49
GF E 42.55 $\pm$ 0.99 32.66 $\pm$ 1.11
F 35.70 $\pm$ 0.78 21.83 $\pm$ 0.27
Table 4: Results for models trained on the ANI-1x data set with ablated weakly supervised losses. All models are trained using 1000 training samples. All results are obtained by averaging over three independent runs. Energy RMSE is given in kcal/mol, while force RMSE is given in kcal/mol/Å.
Model PITC PISC E F
PaiNN 56.62 $\pm$ 2.80 12.96 $\pm$ 0.06
24.60 $\pm$ 0.18 11.51 $\pm$ 0.03
58.30 $\pm$ 2.10 13.18 $\pm$ 0.29
24.53 $\pm$ 0.48 11.43 $\pm$ 0.05
Equiformer 54.52 $\pm$ 4.52 10.10 $\pm$ 0.00
32.64 $\pm$ 26.48 9.64 $\pm$ 0.03
48.96 $\pm$ 4.96 10.30 $\pm$ 0.06
20.89 $\pm$ 0.50 9.68 $\pm$ 0.03
@@
Training Data Set Size Dependence (ANI-1x). We train MLIPs with training set sizes of $[50,10^{2},10^{3},10^{4}]$. The results are plotted in Fig. 2 (a) and (b). Although the observed error reduction depends strongly on the type of MLIP used, the benefit of the weakly supervised losses often decreases slightly with the number of training samples. This result can be expected as the area covered by the weakly supervised losses is also gradually covered by the reference data as the number of training samples increases. Moreover, the gain in accuracy of energy predictions is more significant than that for forces trained only indirectly through the consistency constraint in PITC; see Eq. (3). Finally, it is shown that the improvement is more significant for highly parameterized MLIPs, which benefit the most from increasing the training data size through PIWSL.
@@
Inorganic Bulk Materials (TiO2). Titanium dioxide (TiO2) is a highly relevant metal oxide for industrial applications, featuring several high-pressure phases. Thus, ML models should be able to predict total energies and atomic forces for various high-pressure phases of TiO2, considering periodic boundaries (relevant when aggregating over the local atomic neighborhood). The results for trained models are provided in Table 2. Similar to the ANI-1x data set, our approach improves the accuracy of predicted energies and atomic forces. Interestingly, although the error in the potential energy for 1000 training configurations reaches small RMSE values, from 2 to 4 kcal/mol in predicted energy, the PIWSL still provides a further error reduction. This observation indicates strong evidence of the effectiveness of PIWSL applied to bulk materials.

### v2

@@ 4.2 Physics-Informed Spatial-Consistency Loss
This section introduces a physics-informed approach for generating weak labels based on the concept of conservative forces. Thus, we leverage the fact that the energy difference between two points on the potential energy surface is independent of the path taken between them. We consider two paths from a reference point to the same target point, composed of three perturbation vectors in total. We estimate the potential energy at the target point via Eq. (4.1). An example of two paths is demonstrated in figure 1 (c). The figure relates the energy obtained when displacing atomic positions of the original configuration $\mathcal{S}$ (denoted by A in the figure) by $\delta\mathbf{r}$ (from configuration A to C) with the energy obtained through consecutive perturbations $\delta\mathbf{r}^{\prime}$ (from configuration A to B) and $\delta\mathbf{r}^{\prime\prime}$ (from configuration B to C).
For the first path, we directly predict the energy with an MLIP, i.e., $E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)$, which is related to the approximated energy at $\mathbf{r}+\delta\mathbf{r}$ using Eq. (4.1) through PITC loss. For the second path, we directly compute the energy $E\left(\mathcal{S}_{\delta\mathbf{r}^{\prime}};\bm{\theta}\right)$ for atomic positions displaced by $\delta\mathbf{r}^{\prime}$ and use it to approximate $E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right)$ after applying the second perturbation vector $\delta\mathbf{r}^{\prime\prime}\equiv\delta\mathbf{r}-\delta\mathbf{r}^{\prime}$. The physics-informed spatial consistency (PISC) loss can be defined as
$L_{\mathrm{PISC}}\left(\mathcal{S};\bm{\theta}\right)=\ell\Big{(}E\left(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta}\right),E_{\rm PITC}\left(\mathcal{S}_{\delta\mathbf{r}^{\prime}},\delta\mathbf{r}^{\prime\prime};\bm{\theta}\right)\Big{)},$ (4)
where $E_{\rm PITC}\left(\mathcal{S}_{\delta\mathbf{r}},\delta\mathbf{r}^{\prime};\bm{\theta}\right)$ is the potential energy estimated via PITC formula in Eq. (4.1) from the configuration $\mathcal{S}_{\delta\mathbf{r}}$ perturbed by $\delta\mathbf{r}^{\prime}$. After joint training of PITC and PISC losses, the three different estimations at $\mathcal{S}_{\delta\mathbf{r}}$ become spatially consistent. Note that our conservative forces-based approach is not limited to relations between two perturbation paths or three perturbation vectors. We discuss several other possible configurations in section E.3.
@@ 4.3 Combined Physics-Informed Weakly Supervised Loss
Together with the usual MLIP loss function defined in Eq. (3), the overall objective, which we refer to as the PIWSL loss, can be written as
$\operatorname*{arg\,min}_{\bm{\theta}}\tilde{\mathcal{L}}\left(\mathcal{D};\bm{\theta}\right)=\operatorname*{arg\,min}_{\bm{\theta}}\sum_{\mathcal{S}\in\mathcal{D}}\left(L\left(\mathcal{S};\bm{\theta}\right)\right.$
$\left.+C_{\mathrm{PITC}}L_{\mathrm{PITC}}\left(\mathcal{S};\bm{\theta}\right)+C_{\mathrm{PISC}}L_{\mathrm{PISC}}\left(\mathcal{S};\bm{\theta}\right)\right),$ (5)
@@ 4.4 Determining Perturbation Directions and Magnitudes
The effectiveness of the proposed approach depends on appropriate choices of the perturbation vectors $\delta\mathbf{r}$. We introduce and justify various strategies for generating the perturbations used in Eq. (4.1) and Eq. (4). Any vector $\delta\mathbf{r}$ can be written as $\delta\mathbf{r}\equiv\epsilon\mathbf{g}/\|\mathbf{g}\|_{2}$, where $\epsilon$ is the magnitude of $\delta\mathbf{r}$ and $\mathbf{g}/\|\mathbf{g}\|_{2}$ is its direction. Physical constraints can limit $\epsilon$. Specifically, we can obtain the maximum perturbation length from the validity of the Taylor expansion in Eq. (4.1), which, as discussed in section 5.3, is typically given as at most 30% of the original bond length whose shortest example is the bond between carbon and hydrogen atoms, about 1.09 Å; see also figure 2 (c) and (d). The specific values of $\epsilon$ chosen for our experiments are provided in section C.1.
@@
Second, we compute an adversarial direction, as proposed by Goodfellow et al. (2014); Miyato et al. (2018), which involves defining it as the direction (the gradients) in which the loss error increases the most at the current atom coordinates $\mathbf{r}$ and for the current predicted energy. Assuming the norm of adversarial perturbation as $L_{2}$, the adversarial direction can be approximated by (Miyato et al., 2018)
@@
where $L_{\rm dist}$ is a distance measure function to be maximized by adding $\delta\mathbf{r}_{\rm adv}$, with $\mathbf{y}^{\rm pred}$ and $\mathbf{y}^{\rm ref}$ being the ML model prediction and the reference values. Due to their computational efficiency, we mainly use Eq. (6) in our experiments. A quantitative comparison between the random and adversarial directions is provided in section F.
@@ 5 Experiments
We evaluate our method through extensive experiments with the following objectives: (1) comparing PIWSL with existing baselines, (2) analyzing the impact of the PIWSL using the aspirin molecule, including MD simulations, (3) assessing PIWSL’s ability to improve foundation model finetuning on a sparse dataset, particularly when energy and force predictions when force labels are inaccessible. [^2] In section F, we provide additional results on: (4) a comparison with a prior weakly supervised approach, (5) an ablation study, and (6) a comparison between random and adversarial perturbation vectors. Our experiments mainly use the first-order PITC except for the MD17 and MD22 data sets, where target accuracy requires considering the second-order term. [^3] The effect of the second order term is analyzed in section F.5.
@@ 5.1 Models and Data Sets
We trained the following representative models that are provided in the Open Catalyst code base (Chanussot et al., 2021): SchNet  (Schütt et al., 2017), PaiNN (Schütt et al., 2021), SpinConv (Shuaibi et al., 2021a), eSCN (Passaro & Zitnick, 2023), and Equiformer v2 (Liao et al., 2023), covering MLIPs with a smaller (SchNet, SpinConv, PaiNN) and larger number of parameters (eSCN, Equiformer v2). Moreover, we also considered the MACE model (Batatia et al., 2022), a state-of-the-art model that we use to evaluate the impact of PIWSL on the MD17(CCSD) and MD22 data set. Unless otherwise mentioned and except for SchNet, forces are directly predicted and not computed through the negative gradient of the energy. The results where forces are computed as negative energy gradients are analyzed in section F.1 and section E.10.
To evaluate the effect and dependency of the physics-informed weakly supervised approach in detail, we performed the training on various data sets: ANI-1x as a heterogeneous molecular data set (Smith et al., 2020), TiO2 as a data set for inorganic materials (Artrith & Urban, 2016), the revised MD17 (rMD17) data set containing small molecules with sampled configurational spaces for each (Chmiela et al., 2017, 2018; Christensen & von Lilienfeld, 2020), the MD22 data set containing larger molecules (Chmiela et al., 2023), and LMNTO as another material data set (Cooper et al., 2020); the benchmark results for rMD17, MD22, and LMNTO are provided in section E.1. The detailed description of each data set is provided in section C.3.
@@ 5.2 Benchmark Results
We compare models trained using the PIWSL loss (see Eq. (4.3)) with baseline models trained using the standard supervised loss only (see Eq. (3)). We also compare our approach to a recently proposed data augmentation method that incorporates the task of denoising random perturbations of the atomic coordinates into the learning objective (NoisyNode) (Godwin et al., 2022). A comparison with weak label method (Cooper et al., 2020) is provided in section F. More details on the setup are provided in section C.1. In the following, all evaluation metrics are computed for the test data set.
Heterogeneous Molecular Data Set – ANI-1x. The results provided in Table 1 show that our approach improves the baseline models’ performance in almost all cases. In particular, the error reduction for the predicted energies is often between $10$ and more than $50$. Interestingly, we observe an improved accuracy for potential energies and atomic forces because we include force prediction in PITC and PISC losses, different from the previous work (Cooper et al., 2020). In most cases, except for SchNet, the data augmentation method (NoisyNode) deteriorates the accuracy of the MLIPs because it does not incorporate the proper response of the energy and atomic forces to the perturbation of atomic positions.
@@
$N_{\mathrm{train}}=100$ $N_{\mathrm{train}}=2000$
@@
SchNeta E 17.21 $\pm$ 0.00 19.68 $\pm$ 0.00 17.08 $\pm$ 0.00 10.16 $\pm$ 0.00 $4$$\pm$ 0.00 10.14 $\pm$ 0.00
F 2.84 $\pm$ 0.00 2.70 $\pm$ 0.00 2.83 $\pm$ 0.00 1.87 $\pm$ 0.00 7.45 $\pm$ 0.00 1.85 $\pm$ 0.00
PaiNNb E 14.41 $\pm$ 0.16 n/ab 13.95 $\pm$ 0.09 2.44 $\pm$ 0.03 n/ab 2.30 $\pm$ 0.10
F 1.59 $\pm$ 0.01 n/ab 1.56 $\pm$ 0.01 0.27 $\pm$ 0.02 n/ab 0.24 $\pm$ 0.00
SpinConv E 20.00 $\pm$ 0.42 18.76 $\pm$ 0.74 16.98 $\pm$ 0.99 2.78 $\pm$ 0.67 2.76 $\pm$ 0.42 2.05 $\pm$ 0.39
F 1.58 $\pm$ 0.03 1.53 $\pm$ 0.03 1.59 $\pm$ 0.03 0.67 $\pm$ 0.15 0.61 $\pm$ 0.09 0.61 $\pm$ 0.05
eSCN E 16.41 $\pm$ 1.10 20.92 $\pm$ 0.00 12.63 $\pm$ 0.78 1.78 $\pm$ 0.07 20.90 $\pm$ 0.00 0.90 $\pm$ 0.09
F 1.57 $\pm$ 0.04 1.66 $\pm$ 0.00 1.44 $\pm$ 0.03 0.42 $\pm$ 0.17 1.66 $\pm$ 0.00 0.16 $\pm$ 0.01
Equiformer E 18.21 $\pm$ 0.02 19.06 $\pm$ 0.02 13.93 $\pm$ 0.09 1.73 $\pm$ 0.05 18.54 $\pm$ 0.10 1.27 $\pm$ 0.02
F 1.56 $\pm$ 0.01 1.64 $\pm$ 0.00 1.51 $\pm$ 0.19 0.13 $\pm$ 0.01 1.59 $\pm$ 0.00 0.13 $\pm$ 0.00
@@
a We used a larger batch size of $32$ for SchNet since we obtained extremely high errors for the batch size of $4$. A more detailed discussion of the experimental results for SchNet is provided in section E.1.
@@
b Because of a numerical instability of PaiNN when perturbing atomic coordinates, the cutoff radius is reduced from 12 Å to 5 Å in this experiment. Predicted values become n/a when atomic configurations are perturbed.
@@
Figure 2: (a, b) Relative performance gains for MLIPs trained with PIWSL compared to those trained without it and (c, d) potential energy profiles for a C–H bond of the aspirin molecule. Relative performance gains are evaluated for (a) energy (E-) and (b) force (F-) RMSEs. These results are presented for the ANI-1x data set. Potential energy profiles for a C–H bond of the aspirin molecule are presented for models trained using (c) 100 and (d) 200 configurations. The red and blue arrows indicate the direction from the original structure ($E(\mathcal{S};\bm{\theta})$) to the perturbed one ($E(\mathcal{S}_{\delta\mathbf{r}};\bm{\theta})$), as defined by Eq. (4.1), for the baseline and PIWSL model predictions, respectively.
@@
Training Data Set Size Dependence – ANI-1x. We train MLIPs with training set sizes of $[50,10^{2},10^{3},10^{4},10^{5},$ $10^{6},5\times 10^{6}]$. The results for training data sizes of $10^{5}$, $10^{6}$, and $5\times 10^{6}$ are provided in section E.1. The results are plotted in figure 2 (a) and (b). Although the observed error reduction depends strongly on the type of MLIP used, the benefit of the weakly supervised losses often decreases slightly with the number of training samples. This result can be expected as the area covered by the weakly supervised losses is also gradually covered by the reference data as the number of training samples increases. Moreover, the gain in accuracy of energy predictions is more significant than that for forces trained only indirectly through the consistency constraint in PITC; see Eq. (4.1). Finally, it is shown that the improvement is more significant for highly parameterized MLIPs, which benefit the most from increasing the training data size through PIWSL.
@@
Inorganic Bulk Materials – TiO2. Titanium dioxide (TiO2) is a highly relevant metal oxide for industrial applications, featuring several high-pressure phases. Thus, ML models should be able to predict total energies and atomic forces for various high-pressure phases of TiO2, considering periodic boundaries (relevant when aggregating over the local atomic neighborhood). The results for trained models are provided in Table 2. Similar to the ANI-1x data set, our approach improves the accuracy of predicted energies and atomic forces. Interestingly, even when the error in the potential energy for 2000 training configurations reaches small RMSE values, close to or even less than 1 kcal/mol in predicted energy, the PIWSL still reduces the error. This observation indicates strong evidence of the effectiveness of PIWSL applied to bulk materials.